Mary is 1/4 of Annies's age. joe is ten years old. Jane is 1 and 1/3 times Mary's age. Joe is 2/3 of Mary's age. What are all of their ages? Who has great grand children?
Answer by stanbon(75887)

Mary is 1/4 of Annies's age. joe is ten years old. Jane is 1 and 1/3 times Mary's age. Joe is 2/3 of Mary's age. What are all of their ages? Who has great grand children?
Equations:
m = (1/4)a
j = 10
ja = (4/3)m
j = (2/3)m
-----------
If joe is 10 yrs, mary is (3/2)10 = 15 yrs
Then Annie is 4*15 = 60 yrs and Jane = (4/3)15 = 20 yrs
-------
Annie may have great grandchildren.
Cheers,
Stan H.
------------